# Distributionally Robust Beamforming and Estimation of Wireless Signals

### v1

@@ III-C Distibutionally Robust Estimation
where $\mathcal{U}_{\mathbf{x},\mathbf{s}}$, called a distributional uncertainty set, contains a collection of distributions that are close to the nominal distribution (i.e., the sample-estimated distribution) ^$\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}}$, i.e.,
@@
where $d(\cdot,\cdot)$ denotes a distance between two distributions. Since ^$\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}}$ is discrete and $\mathbb{P}_{\mathbf{x},\mathbf{s}}$ can be continuous, the Wasserstein distance [27, Def. 2] and the maximum mean discrepancy (MMD) distance [28, Def. 2.1] are the usual choices of $d(\cdot,\cdot)$ to construct $\mathcal{U}_{\mathbf{x},\mathbf{s}}$. For extensive reading on the Wasserstein and MMD distances, see Appendix D of the online supplementary materials. When $\epsilon=0$, (17) reduces to (15). If $\mathcal{U}_{\mathbf{x},\mathbf{s}}$ contains only Gaussian distributions, (17) is particularized to
$\begin{array}[]{cl}\displaystyle\min_{\bm{W}}\max_{\bm{R}_{x},\bm{R}_{xs},\bm{R}_{s}}&\operatorname{Tr}\big{[}\bm{W}\bm{R}_{x}\bm{W}^{\mathsf{H}}-\bm{W}\bm{R}_{xs}-\bm{R}^{\mathsf{H}}_{xs}\bm{W}^{\mathsf{H}}+\bm{R}_{s}\big{]}\\ \text{s.t.}&d_{1}(\bm{R}_{x},{\hat{\bm{R}}}_{x})\leq\epsilon_{1},\\ &d_{2}(\bm{R}_{xs},{\hat{\bm{R}}}_{xs})\leq\epsilon_{2},\\ &d_{3}(\bm{R}_{s},{\hat{\bm{R}}}_{s})\leq\epsilon_{3},\\ &\bm{R}_{x}\succeq\bm{0},~{}\bm{R}_{x}=\bm{R}^{\mathsf{H}}_{x},\\ &\bm{R}_{s}\succeq\bm{0},~{}\bm{R}_{s}=\bm{R}^{\mathsf{H}}_{s},\end{array}$ (19)
because every zero-mean complex Gaussian distribution is uniquely characterized by its covariance and pseudo-covariance, but in beamforming, we do not consider pseudo-covariances; cf. (13); $d_{1}$, $d_{2}$, and $d_{3}$ are matrix similarity measures (e.g., matrix distances); $\epsilon_{1}\geq 0$, $\epsilon_{2}\geq 0$, and $\epsilon_{3}\geq 0$ are uncertainty quantification parameters. When $\epsilon_{1}=\epsilon_{2}=\epsilon_{3}=0$, (19) reduces to (16). The fact below characterizes the performance of the distributionally robust estimators.
Fact 1
Suppose that the true distribution $\mathbb{P}_{0,\mathbf{x},\mathbf{s}}$ of $(\mathbf{x},\mathbf{s})$ is included in $\mathcal{U}_{\mathbf{x},\mathbf{s}}$; for notational clarity, we hereafter distinguish $\mathbb{P}_{0,\mathbf{x},\mathbf{s}}$ from $\mathbb{P}_{\mathbf{x},\mathbf{s}}$. The true estimation error evaluated at $\mathbb{P}_{0,\mathbf{x},\mathbf{s}}$
$\min_{\bm{\phi}\in\mathcal{B}}\operatorname{Tr}\mathbb{E}_{(\mathbf{x},\mathbf{s})\sim\mathbb{P}_{0,\mathbf{x},\mathbf{s}}}[\bm{\phi}(\mathbf{x})-\mathbf{s}][\bm{\phi}(\mathbf{x})-\mathbf{s}]^{\mathsf{H}}$ (20)
is upper bounded by the worst-case estimation error (17). Therefore, by diminishing the upper bound in (17), the true estimation error (20) can also be reduced. However, the conventional empirical estimation error (15) cannot upper bound the true estimation error (20). This performance guarantee is the benefit of considering the distributionally robust method (17). Due to the weak convergence property of the empirical distribution to the true data-generating distribution, that is, $d(\mathbb{P}_{0,\mathbf{x},\mathbf{s}},~{}\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}})\to 0$ as the sample size $L\to\infty$, there exists $\epsilon$ in (18) for every $L$, such that $\mathbb{P}_{0,\mathbf{x},\mathbf{s}}$ is included in $\mathcal{U}_{\mathbf{x},\mathbf{s}}$ in $\mathbb{P}^{L}_{0,\mathbf{x},\mathbf{s}}$-probability ($L$-fold product measure of $\mathbb{P}_{0,\mathbf{x},\mathbf{s}}$). normal-$\square$
@@ IV-A General Framework and Concrete Examples
The following lemma equivalently transforms Problem (19).
@@
Problem (19) is equivalent to
$\begin{array}[]{cl}\displaystyle\max_{\bm{R}_{x},\bm{R}_{xs},\bm{R}_{s}}\min_{\bm{W}}&\operatorname{Tr}\big{[}\bm{W}\bm{R}_{x}\bm{W}^{\mathsf{H}}-\bm{W}\bm{R}_{xs}-\bm{R}^{\mathsf{H}}_{xs}\bm{W}^{\mathsf{H}}+\bm{R}_{s}\big{]}\\ \text{s.t.}&d_{1}(\bm{R}_{x},{\hat{\bm{R}}}_{x})\leq\epsilon_{1},\\ &d_{2}(\bm{R}_{xs},{\hat{\bm{R}}}_{xs})\leq\epsilon_{2},\\ &d_{3}(\bm{R}_{s},{\hat{\bm{R}}}_{s})\leq\epsilon_{3},\\ &\bm{R}_{x}\succeq\bm{0},~{}\bm{R}_{x}=\bm{R}^{\mathsf{H}}_{x},\\ &\bm{R}_{s}\succeq\bm{0},~{}\bm{R}_{s}=\bm{R}^{\mathsf{H}}_{s},\end{array}$ (21)
provided that $d_{1}(\bm{R}_{x},{\hat{\bm{R}}}_{x})$ is convex in $\bm{R}_{x}$, $d_{2}(\bm{R}_{xs},{\hat{\bm{R}}}_{xs})$ is convex in $\bm{R}_{xs}$, and $d_{3}(\bm{R}_{s},{\hat{\bm{R}}}_{s})$ is convex in $\bm{R}_{s}$. Problem (19) is further equivalent to
$\begin{array}[]{cl}\displaystyle\max_{\bm{R}_{x},\bm{R}_{xs},\bm{R}_{s}}&\operatorname{Tr}\big{[}-\bm{R}^{\mathsf{H}}_{xs}\bm{R}^{-1}_{x}\bm{R}_{xs}+\bm{R}_{s}\big{]}\\ \text{s.t.}&d_{1}(\bm{R}_{x},{\hat{\bm{R}}}_{x})\leq\epsilon_{1},\\ &d_{2}(\bm{R}_{xs},{\hat{\bm{R}}}_{xs})\leq\epsilon_{2},\\ &d_{3}(\bm{R}_{s},{\hat{\bm{R}}}_{s})\leq\epsilon_{3},\\ &\bm{R}_{x}\succeq\bm{0},~{}\bm{R}_{x}=\bm{R}^{\mathsf{H}}_{x},\\ &\bm{R}_{s}\succeq\bm{0},~{}\bm{R}_{s}=\bm{R}^{\mathsf{H}}_{s}.\end{array}$ (22)
@@
This is by the Sion’s minimax theorem. Note that the objective function is convex in $\bm{W}$ and linear (thus concave) in both $\bm{R}_{x}$, $\bm{R}_{xs}$, and $\bm{R}_{s}$. Note also that complex Euclidean spaces are linear topological spaces. For every given $\bm{R}_{x}$ and $\bm{R}_{xs}$, the inner minimization sub-problem of (21) is solved by the Wiener beamformer $\bm{W}^{\star}_{\text{Wiener}}=\bm{R}^{\mathsf{H}}_{xs}\bm{R}^{-1}_{x}$, which transforms (21) to 22. This completes the proof. $\square$
@@
$f(\bm{R}_{x})\coloneqq\operatorname{Tr}\big{[}-\bm{R}^{\mathsf{H}}_{xs}\bm{R}^{-1}_{x}\bm{R}_{xs}+\bm{R}_{s}\big{]}$ (23)
denote the objective function of (22) in terms of $\bm{R}_{x}$, i.e., a matrix-valued function of $\bm{R}_{x}\succeq\bm{0}$. The following theorem shows the monotonicity of $f(\bm{R}_{x})$ in $\bm{R}_{x}$.
@@
Function $f(\bm{R}_{x})$ in (23) is monotonically increasing in $\bm{R}_{x}$: i.e., if $\bm{R}_{x,1}\succeq\bm{R}_{x,2}$, we have $f(\bm{R}_{x,1})\geq f(\bm{R}_{x,2})$.
@@
Theorem 1 implies that if $\bm{R}^{\star}_{x}\succeq\bm{R}_{x}$ for all $\bm{R}_{x}$ such that $d_{1}(\bm{R}_{x},{\hat{\bm{R}}}_{x})\leq\epsilon_{1}$, then $\bm{R}^{\star}_{x}$ is a maximizer of Problem (22).
It is also straightforward to see that if $\bm{R}^{\star}_{s}\succeq\bm{R}_{s}$ for all $\bm{R}_{s}$ such that $d_{3}(\bm{R}_{s},{\hat{\bm{R}}}_{s})\leq\epsilon_{3}$, then $\bm{R}^{\star}_{s}$ is a maximizer of Problem (22).
The objective function of Problem (22) is negative-definite concave in $\operatorname{vec}(\bm{R}_{xs})$. Therefore, Problem (22) can be readily solved using any first-order method such as the projected gradient ascent method; the gradient of the objective function with respect to $\bm{R}_{xs}$ is $-2\bm{R}^{-1}_{x}\bm{R}_{xs}$. When we do not study the uncertainty in the channel $\bm{H}$ (because this paper does not intend to explicitly estimate the channel matrix), since $\bm{R}_{xs}=\bm{H}\bm{P}$, the uncertainty of $\bm{R}_{xs}$ is induced by the uncertainty of the covariance $\bm{P}\coloneqq\bm{R}_{s}$ of the transmitted signal $\mathbf{s}$. As a result, we can study the uncertainty in $\bm{P}$ instead.
If the uncertainties in $\bm{R}_{x}$ and $\bm{R}_{xs}$ are induced by the uncertainties in the transmitted signal covariance $\bm{P}\coloneqq\bm{R}_{s}$ and the noise covariance $\bm{R}\coloneqq\bm{R}_{v}$, Problem (22) can be explicitly written as
$\begin{array}[]{cl}\displaystyle\max_{\bm{P},\bm{R}}&\operatorname{Tr}\big{[}\bm{P}-\bm{P}\bm{H}^{\mathsf{H}}(\bm{H}\bm{P}\bm{H}^{\mathsf{H}}+\bm{R})^{-1}\bm{H}\bm{P}\big{]}\\ \text{s.t.}&d_{4}(\bm{P},\hat{\bm{P}})\leq\epsilon_{4},\\ &d_{5}(\bm{R},\hat{\bm{R}})\leq\epsilon_{5},\\ &\bm{P}\succeq\bm{0},~{}\bm{R}\succeq\bm{0},\\ &\bm{P}=\bm{P}^{\mathsf{H}},~{}\bm{R}=\bm{R}^{\mathsf{H}},\end{array}$ (24)
for some similarity measures $d_{4}$ and $d_{5}$ and nonnegative scalars $\epsilon_{4}$ and $\epsilon_{5}$. For every $(\bm{P},\bm{R})$, the associated beamformer is given in (13). When the uncertainty in the channel matrix must be investigated, we can consider
$\begin{array}[]{cl}\displaystyle\max_{\bm{H}}&\operatorname{Tr}\big{[}\bm{P}-\bm{P}\bm{H}^{\mathsf{H}}(\bm{H}\bm{P}\bm{H}^{\mathsf{H}}+\bm{R})^{-1}\bm{H}\bm{P}\big{]}\\ \text{s.t.}&d_{6}(\bm{H},\hat{\bm{H}})\leq\epsilon_{6},\end{array}$ (25)
which is not a semi-definite program. In addition, the gradient of the objective function with respect to $\bm{H}$ is complicated to obtain. Hence, a solution method should be tailored.
Below we provide some concrete examples to showcase the practical usefulness and applications of (22) and (24).
Example 1 (Diagonally Loaded Beamforming)
Consider the uncertainty in $\bm{R}_{x}$. Suppose $d_{1}(\bm{R}_{x},{\hat{\bm{R}}}_{x})\leq\epsilon_{1}$ is particularized as
${\hat{\bm{R}}}_{x}-\epsilon_{1}\bm{I}_{N}\preceq\bm{R}_{x}\preceq{\hat{\bm{R}}}_{x}+\epsilon_{1}\bm{I}_{N};$ (26)
that is, $\bm{R}_{x}$ lies in the diagonal $\epsilon_{1}$-neighbourhood (i.e., diagonal $\epsilon_{1}$-perturbation) of ${\hat{\bm{R}}}_{x}$. By Theorem 1, we have the distributionally robust beamformer
@@
$\begin{array}[]{cl}\bm{W}^{\star}_{\text{DR-DL}}&=\bm{R}^{\mathsf{H}}_{xs}[{\hat{\bm{R}}}_{x}+\epsilon_{1}\bm{I}_{N}]^{-1}\\ &=\bm{P}\bm{H}^{\mathsf{H}}[\bm{H}\bm{P}\bm{H}^{\mathsf{H}}+\hat{\bm{R}}+\epsilon_{1}\bm{I}_{N}]^{-1},\end{array}$ (27)
which is known as the loaded sample matrix inversion method or the diagonal loading method [19], [14, Eq. (11)]. Note that when $\bm{R}_{xs}$ is accurate, $\bm{P}$ and $\bm{H}$ are accurate. Therefore, the uncertainty in $\hat{\bm{R}}_{x}$ comes from the uncertainty in $\hat{\bm{R}}$. normal-$\square$
Diagonally loaded beamforming can handle the scarcity of pilot data: When $L$ is not sufficiently large, the estimate $\hat{\bm{R}}_{x}$ of $\bm{R}_{x}$ would be significantly unreliable; cf. Subsection III-B.

### v2

@@ III-B2 Distibutionally Robust Estimation
where $\mathcal{U}_{\mathbf{x},\mathbf{s}}$, called a distributional uncertainty set, contains a collection of distributions that are close to the nominal distribution (i.e., the sample-estimated distribution) ^$\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}}$;
@@
where $d(\cdot,\cdot)$ denotes a similarity measure (e.g., metric or divergence) between two distributions and $\epsilon\geq 0$ an uncertainty quantification level. Since ^$\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}}$ is discrete and $\mathbb{P}_{\mathbf{x},\mathbf{s}}$ is not, the Wasserstein distance [27, Def. 2] and the maximum mean discrepancy (MMD) distance [28, Def. 2.1] are the typical choices of $d(\cdot,\cdot)$ to construct $\mathcal{U}_{\mathbf{x},\mathbf{s}}$. When ^$\hat{\mathbb{P}}_{\mathbf{x},\mathbf{s}}$ and $\mathbb{P}_{\mathbf{x},\mathbf{s}}$ are parametric distributions (e.g., Gaussian, exponential family), divergences such as the Kullback-–Leibler (KL) divergence, or more general $\phi$-divergence, are also applicable to particularize $d(\cdot,\cdot)$ because parameters can be estimated using samples. When $\epsilon=0$, (17) reduces to (15).
If $\mathcal{U}_{\mathbf{x},\mathbf{s}}$ contains only Gaussian distributions, (17) is particularized to
$\begin{array}[]{cl}\displaystyle\min_{\bm{W}}\max_{\bm{R}}&\operatorname{Tr}\big{[}\bm{W}\bm{R}_{x}\bm{W}^{\mathsf{H}}-\bm{W}\bm{R}_{xs}-\bm{R}^{\mathsf{H}}_{xs}\bm{W}^{\mathsf{H}}+\bm{R}_{s}\big{]}\\ \text{s.t.}&{d_{0}(\bm{R},~{}\hat{\bm{R}})\leq\epsilon_{0}},\\ &{\bm{R}\succeq\bm{0},}\end{array}$ (19)
where
${\bm{R}\coloneqq\left[\begin{array}[]{cc}\bm{R}_{x}&\bm{R}_{xs}\\ \bm{R}^{\mathsf{H}}_{xs}&\bm{R}_{s}\end{array}\right],}~{}~{}~{}\hat{\bm{R}}\coloneqq\left[\begin{array}[]{cc}\hat{\bm{R}}_{x}&\hat{\bm{R}}_{xs}\\ \hat{\bm{R}}^{\mathsf{H}}_{xs}&\hat{\bm{R}}_{s}\end{array}\right],$ (20)
because every zero-mean complex Gaussian distribution is uniquely characterized by its covariance and pseudo-covariance, but in receive beamforming, we do not consider pseudo-covariances; cf. (13); $d_{0}$ denotes the matrix similarity measures (e.g., matrix distances); $\epsilon_{0}\geq 0$ is the uncertainty quantification parameter. When $\epsilon_{0}=0$, (19) reduces to (16).
@@ IV-A General Framework and Concrete Examples
The following lemma solves Problem (19).
@@
Suppose that the constraint $d_{0}(\bm{R},~{}\hat{\bm{R}})\leq\epsilon_{0}$ is compact convex and $\bm{R}_{x}$ is invertible. Let $\bm{R}^{\star}$ solve the problem below:
$\begin{array}[]{cl}\displaystyle\max_{\bm{R}}&\operatorname{Tr}\big{[}-\bm{R}^{\mathsf{H}}_{xs}\bm{R}^{-1}_{x}\bm{R}_{xs}+\bm{R}_{s}\big{]}\\ \text{s.t.}&d_{0}(\bm{R},~{}\hat{\bm{R}})\leq\epsilon_{0},\\ &\bm{R}\succeq\bm{0},~{}~{}~{}\bm{R}_{x}\succ\bm{0}.\end{array}$ (21)
Construct $\bm{W}^{\star}$ using $\bm{R}^{\star}$ as follows:
$\bm{W}^{\star}\coloneqq\bm{R}^{\star\mathsf{H}}_{xs}\bm{R}^{\star-1}_{x}.$ (22)
Then $(\bm{W}^{\star},\bm{R}^{\star})$ is a solution to Problem (19). On the other hand, if $(\bm{W}^{\prime},\bm{R}^{\prime})$ solves Problem (19), then $\bm{R}^{\prime}$ is a solution to (21) and $(\bm{W}^{\prime},\bm{R}^{\prime})$ satisfies (22).
@@
See Appendix D of the online supplementary materials. $\square$
@@
$f_{1}(\bm{R})\coloneqq\operatorname{Tr}\big{[}-\bm{R}^{\mathsf{H}}_{xs}\bm{R}^{-1}_{x}\bm{R}_{xs}+\bm{R}_{s}\big{]}$ (23)
denote the objective function of (21). When $\bm{R}_{s}$ and $\bm{R}_{xs}$ are fixed, we define
$f_{2}(\bm{R}_{x})\coloneqq\operatorname{Tr}\big{[}-\bm{R}^{\mathsf{H}}_{xs}\bm{R}^{-1}_{x}\bm{R}_{xs}+\bm{R}_{s}\big{]}.$ (24)
The theorem below studies the properties of $f_{1}$ and $f_{2}$.
@@
Consider the definition of $\bm{R}$ in (20). The functions $f_{1}$ defined in (23) and $f_{2}$ defined in (24) are monotonically increasing in $\bm{R}$ and $\bm{R}_{x}$, respectively. To be specific, if $\bm{R}_{1}\succeq\bm{R}_{2}\succeq\bm{0}$, $\bm{R}_{1,x}\succ\bm{0}$, and $\bm{R}_{2,x}\succ\bm{0}$, we have $f_{1}(\bm{R}_{1})\geq f_{1}(\bm{R}_{2})$. In addition, if $\bm{R}_{1,x}\succeq\bm{R}_{2,x}\succ\bm{0}$, we have $f_{2}(\bm{R}_{1,x})\geq f_{2}(\bm{R}_{2,x})$.
@@
To concretely solve (21), we need to particularize $d_{0}$. This article investigates the following uncertainty sets.
Definition 1 (Additive Moment Uncertainty Set)
The additive moment uncertainty set of $\bm{R}$ is constructed as
$\hat{\bm{R}}-\epsilon_{0}\bm{E}\preceq\bm{R}\preceq\hat{\bm{R}}+\epsilon_{0}\bm{E},$ (25)
for some $\bm{E}\succeq\bm{0}$ and $\epsilon_{0}\geq 0$ such that $\hat{\bm{R}}-\epsilon_{0}\bm{E}\succeq\bm{0}$. $\square$
Definition 1 is motivated by the fact that the difference $\bm{R}-\hat{\bm{R}}$ is bounded by some threshold matrix $\bm{E}$ and error quantification level $\epsilon_{0}$: specifically, $-\epsilon_{0}\bm{E}\preceq\bm{R}-\hat{\bm{R}}\preceq\epsilon_{0}\bm{E}$. In practice, we can consider the threshold as an identity matrix because, for every non-identity $\bm{E}\succeq\bm{0}$, we have $\bm{E}\preceq\lambda_{1}\bm{I}_{N+M}$ where $\lambda_{1}$ is the largest eigenvalue of $\bm{E}$.
Definition 2 (Diagonal-Loading Uncertainty Set)
The diagonal-loading uncertainty set of $\bm{R}$ is constructed as
$\hat{\bm{R}}-\epsilon_{0}\bm{I}_{N+M}\preceq\bm{R}\preceq\hat{\bm{R}}+\epsilon_{0}\bm{I}_{N+M},$ (26)
for some $\epsilon_{0}\geq 0$ such that $\hat{\bm{R}}-\epsilon_{0}\bm{I}_{N+M}\succeq\bm{0}$. $\square$
Due to the concentration property of the sample-covariance $\hat{\bm{R}}$ to the true covariance $\bm{R}$ when the true distribution $\mathbb{P}_{\mathbf{x},\mathbf{s}}$ is fixed within a frame, finite values of $\epsilon_{0}$ exist for every sample size $L$; NB: $\epsilon_{0}\to 0$ as $L\to\infty$. However, given $L$, the smallest $\epsilon_{0}$ cannot be practically calculated because it depends on the true but unknown $\mathbb{P}_{\mathbf{x},\mathbf{s}}$. If $\bm{E}$ is block-diagonal, the generalized diagonal-loading uncertainty set can be motivated.
Definition 3 (Generalized Diagonal-Loading Uncertainty Set)
The generalized diagonal-loading uncertainty set of $\bm{R}$ is constructed as
@@
$\begin{array}[]{l}\left[\begin{array}[]{cc}\hat{\bm{R}}_{x}&\hat{\bm{R}}_{xs}\\ \hat{\bm{R}}^{\mathsf{H}}_{xs}&\hat{\bm{R}}_{s}\end{array}\right]-\epsilon_{0}\left[\begin{array}[]{cc}\bm{F}&\bm{0}\\ \bm{0}&\bm{G}\end{array}\right]\\ \quad\quad\preceq\left[\begin{array}[]{cc}\bm{R}_{x}&\bm{R}_{xs}\\ \bm{R}^{\mathsf{H}}_{xs}&\bm{R}_{s}\end{array}\right]\\ \quad\quad\quad\quad\preceq\left[\begin{array}[]{cc}\hat{\bm{R}}_{x}&\hat{\bm{R}}_{xs}\\ \hat{\bm{R}}^{\mathsf{H}}_{xs}&\hat{\bm{R}}_{s}\end{array}\right]+\epsilon_{0}\left[\begin{array}[]{cc}\bm{F}&\bm{0}\\ \bm{0}&\bm{G}\end{array}\right],\end{array}$ (27)
for some $\bm{F},\bm{G}\succeq\bm{0}$ and $\epsilon_{0}\geq 0$ such that the matrix in the first line is positive semi-definite. $\square$
Definitions 1, 2, and 3 are introduced for the first time in this article. Another type of moment-based uncertainty set is popular in the literature, which we refer to as the multiplicative moment uncertainty set for differentiation.
Definition 4 (Multiplicative Moment Uncertainty Set [29])
The multiplicative moment uncertainty set of $\bm{R}$ is given as
$\theta_{1}\hat{\bm{R}}\preceq\bm{R}\preceq\theta_{2}\hat{\bm{R}},$ (28)
for some $\theta_{2}\geq 1\geq\theta_{1}\geq 0$. $\square$
The following corollary shows the distributionally robust linear beamformers associated with the various uncertainty sets in Definitions 1, 2, 3, and 4.
Corollary 1 (of Theorem 1)
Consider the moment-based uncertainty sets in Definitions 1, 2, 3, and 4. The distributionally robust linear beamforming (21) is analytically solved by the corresponding upper bounds of $\bm{R}$. To be specific,
  1.  C1)
    Under Definition 1, the additive-moment distributionally robust (DR-AM) beamformer is
    $\begin{array}[]{cl}\bm{W}^{\star}_{\text{DR-AM}}&=(\hat{\bm{R}}_{xs}+\epsilon_{0}\bm{E}_{xs})^{\mathsf{H}}(\hat{\bm{R}}_{x}+\epsilon_{0}\bm{E}_{x})^{-1}\\ &=(\hat{\bm{H}}\hat{\bm{R}}_{s}+\epsilon_{0}\bm{E}_{xs})^{\mathsf{H}}\cdot\\ &\quad\quad\quad[\hat{\bm{H}}\hat{\bm{R}}_{s}\hat{\bm{H}}^{\mathsf{H}}+\hat{\bm{R}}_{v}+\epsilon_{0}\bm{E}_{x}]^{-1}.\end{array}$ (29)
  2.  C2)
    Under Definition 2, the diagnal-loading distributionally robust (DR-DL) beamformer is
    $\begin{array}[]{cl}\bm{W}^{\star}_{\text{DR-DL}}&=\hat{\bm{R}}^{\mathsf{H}}_{xs}[{\hat{\bm{R}}}_{x}+\epsilon_{0}\bm{I}_{N}]^{-1}\\ &=\hat{\bm{R}}_{s}\hat{\bm{H}}^{\mathsf{H}}[\hat{\bm{H}}\hat{\bm{R}}_{s}\hat{\bm{H}}^{\mathsf{H}}+\hat{\bm{R}}_{v}+\epsilon_{0}\bm{I}_{N}]^{-1},\end{array}$ (30)
    which is also known as the loaded sample matrix inversion method [19], [14, Eq. (11)] and widely-used in the practice of wireless communications.
  3.  C3)
    Under Definition 3, the generalized diagonal-loading distributionally robust beamformer (DR-GDL) is
    $\begin{array}[]{cl}\bm{W}^{\star}_{\text{DR-GDL}}&=\hat{\bm{R}}^{\mathsf{H}}_{xs}[{\hat{\bm{R}}}_{x}+\epsilon_{0}\bm{F}]^{-1}\\ &=\hat{\bm{R}}_{s}\hat{\bm{H}}^{\mathsf{H}}[\hat{\bm{H}}\hat{\bm{R}}_{s}\hat{\bm{H}}^{\mathsf{H}}+\hat{\bm{R}}_{v}+\epsilon_{0}\bm{F}]^{-1}.\end{array}$ (31)
  4.  C4)
    Under Definition 4, the multiplicative-moment (MM) distributionally robust beamformer is identical to the Wiener beamformer (13) at nominal values:
    $\begin{array}[]{cl}\bm{W}^{\star}_{\text{DR-MM}}&=\hat{\bm{R}}_{xs}^{\mathsf{H}}\hat{\bm{R}}_{x}^{-1}\\ &=\hat{\bm{R}}_{s}\hat{\bm{H}}^{\mathsf{H}}[\hat{\bm{H}}\hat{\bm{R}}_{s}\hat{\bm{H}}^{\mathsf{H}}+\hat{\bm{R}}_{v}]^{-1}.\end{array}$ (32)
The corresponding estimation errors are simple to obtain. $\square$
Corollary 1 implies that, in the sense of the same induced robust beamformers, the diagonal-loading uncertainty set (26) and the generalized diagonal-loading uncertainty set (27) are technically equivalent to the following trimmed versions.
Definition 5 (Trimmed Diagonal-Loading Uncertainty Sets)
By setting $\bm{G}\coloneqq\bm{0}$ in (27), in terms of $\bm{R}_{x}$, (27) reduces to the trimmed generalized diagonal-loading uncertainty set:
$\hat{\bm{R}}_{x}-\epsilon_{0}\bm{F}\preceq\bm{R}_{x}\preceq\hat{\bm{R}}_{x}+\epsilon_{0}\bm{F}.$ (33)
The trimmed diagonal-loading uncertainty set
$\hat{\bm{R}}_{x}-\epsilon_{0}\bm{I}_{N}\preceq\bm{R}_{x}\preceq\hat{\bm{R}}_{x}+\epsilon_{0}\bm{I}_{N},$ (34)
is obtained by letting $\bm{F}\coloneqq\bm{I}_{N}$. $\square$
The robust beamformers corresponding to the trimmed uncertainty sets (33) and (34) remain the same as defined in (31) and (30), respectively; cf. Theorem 1.
As we can see from Corollary 1, the primary benefit of using the moment-based uncertainty sets is the computational simplicity due to the availability of closed-form solutions. If the uncertainty sets are constructed using the Wasserstein distance $\sqrt{\operatorname{Tr}[{\bm{R}+\hat{\bm{R}}-2(\hat{\bm{R}}^{1/2}\bm{R}\hat{\bm{R}}^{1/2})^{1/2}}]}\leq\epsilon_{0}$ or the KL divergence $\frac{1}{2}[\operatorname{Tr}[{\hat{\bm{R}}^{-1}\bm{R}-\bm{I}_{N+M}}]-\ln\det(\hat{\bm{R}}^{-1}\bm{R})]\leq\epsilon_{0}$ between $\mathcal{CN}(\bm{0},\bm{R})$ and $\mathcal{CN}(\bm{0},\hat{\bm{R}})$, the induced distributionally robust linear beamforming problems have no closed-form solutions, and therefore, are computationally prohibitive in practice. In addition, Corollary 1 suggests that the distributionally robust beamformer under the multiplicative moment uncertainty set (28) is the same as the nominal beamformer $\hat{\bm{R}}^{\mathsf{H}}_{xs}\hat{\bm{R}}_{x}^{-1}$, which essentially do not introduce robustness in wireless signal estimation; this is another motivation why we construct new moment-based uncertainty sets in Definitions 1, 2, and 3. However, we can modify the multiplicative moment uncertainty set in Definition 4 to achieve robustness.
Definition 6 (Modified Multiplicative Moment Uncertainty Set)
The modified multiplicative moment uncertainty set of $\bm{R}$ is given as
$\left[\begin{array}[]{cc}\theta_{1}\hat{\bm{R}}_{x}&\hat{\bm{R}}_{xs}\\ \hat{\bm{R}}^{\mathsf{H}}_{xs}&\theta_{1}\hat{\bm{R}}_{s}\end{array}\right]\preceq\left[\begin{array}[]{cc}\bm{R}_{x}&\bm{R}_{xs}\\ \bm{R}^{\mathsf{H}}_{xs}&\bm{R}_{s}\end{array}\right]\preceq\left[\begin{array}[]{cc}\theta_{2}\hat{\bm{R}}_{x}&\hat{\bm{R}}_{xs}\\ \hat{\bm{R}}^{\mathsf{H}}_{xs}&\theta_{2}\hat{\bm{R}}_{s}\end{array}\right]$ (35)
for some $\theta_{2}\geq 1\geq\theta_{1}\geq 0$ such that the left-most matrix is positive semi-definite. $\square$
The robust beamformer under the modified multiplicative moment uncertainty set (35) is $\bm{W}^{\star}_{\text{DR-MMM}}=\hat{\bm{R}}^{\mathsf{H}}_{xs}\cdot[\theta_{2}\hat{\bm{R}}_{x}]^{-1}$.
In terms of the uncertainties of $\bm{R}_{s}$ and $\bm{R}_{v}$, Problem (21) can be explicitly written as
$\begin{array}[]{cl}\displaystyle\max_{\bm{R}_{s},\bm{R}_{v}}&\operatorname{Tr}\big{[}\bm{R}_{s}-\bm{R}_{s}\bm{H}^{\mathsf{H}}(\bm{H}\bm{R}_{s}\bm{H}^{\mathsf{H}}+\bm{R}_{v})^{-1}\bm{H}\bm{R}_{s}\big{]}\\ \text{s.t.}&d_{1}(\bm{R}_{s},\hat{\bm{R}}_{s})\leq\epsilon_{1},\\ &d_{2}(\bm{R}_{v},\hat{\bm{R}}_{v})\leq\epsilon_{2},\\ &\bm{R}_{s}\succeq\bm{0},~{}\bm{R}_{v}\succeq\bm{0},\end{array}$ (36)
for some similarity measures $d_{1}$ and $d_{2}$ and nonnegative scalars $\epsilon_{1}$ and $\epsilon_{2}$. For every given $(\bm{R}_{s},\bm{R}_{v})$, the associated beamformer is given in (13). When the uncertainty in the channel matrix must be investigated, we can consider
$\begin{array}[]{cl}\displaystyle\max_{\bm{H}}&\operatorname{Tr}\big{[}\bm{R}_{s}-\bm{R}_{s}\bm{H}^{\mathsf{H}}(\bm{H}\bm{R}_{s}\bm{H}^{\mathsf{H}}+\bm{R}_{v})^{-1}\bm{H}\bm{R}_{s}\big{]}\\ \text{s.t.}&d_{3}(\bm{H},\hat{\bm{H}})\leq\epsilon_{3},\end{array}$ (37)
which is not a semi-definite program. In addition, the gradient of the objective function with respect to $\bm{H}$ is complicated to obtain. Hence, practically, we should avoid directly attacking Problem (37); this can be done by directly considering the uncertainties of $\bm{R}_{x}$ and $\bm{R}_{xs}$ (i.e., $\bm{R}$) because the uncertainties of $\bm{R}_{s}$, $\bm{R}_{v}$, and $\bm{H}$ can be reflected in the uncertainties of $\bm{R}_{x}$ and $\bm{R}_{xs}$; cf. $\bm{R}_{x}=\bm{H}\bm{R}_{s}\bm{H}^{\mathsf{H}}+\bm{R}_{v}$ and $\bm{R}_{xs}=\bm{H}\bm{R}_{s}$.
In addition to Corollary 1, below we provide other concrete examples to further showcase the usefulness and applications of the distributionally robust beamforming formulations (21) and (36), where the trimmed uncertainty sets are employed.